# Non-conformally Einstein instantons in Conformal Gravity with and without nonlinear matter fields

### v1

@@ Abstract
In this work, we study non-conformally Einstein gravitational instantons in four-dimensional Conformal Gravity, both in vacuum and in the presence of nonlinear conformal matter. First, the one-parameter extension of the Kerr-NUT-AdS metric is analyzed. We obtain their conserved charges by using the Noether-Wald formalism. It turns out that they receive corrections from the linear modes present in Conformal Gravity, which are properly identified. Then, we perform the analytic continuation into the Euclidean section and find the curve in parameter space along which this solution becomes regular and globally (anti)-self-dual. Using the Dunajski-Tod theorem, we show that the metric is not conformally Einstein. Then, the backreaction of nonlinear conformal matter is considered. In particular, we find new gravitational instantons in the presence of conformally coupled scalar fields and ModMax electrodynamics. We compute the partition function and conserved charges, which turn out to be finite by virtue of the conformal invariance of the theory. As a byproduct, we also obtain a generalization of the Riegert metric dressed with nonlinear conformal matter as a particular limit of these instantons. For all cases, we analyze the global properties, the curve in parameter space where the solutions are (anti)-self-dual, and the on-shell Euclidean action, among other features.
@@ I Introduction
The AdS/CFT correspondence has provided a powerful tool for studying strongly interacting field theories at the fixed points of the renormalization group flow, where standard perturbative techniques are no longer valid Maldacena (1998); Witten (1998); Gubser et al. (1998). At those points, quantum field theories exhibit conformal invariance, which dictates the specific form of correlator functions of the theory. In the context of gravity, theories that possess conformal symmetry display an improved ultraviolet behavior, providing a suitable scenario for studying quantum effects. Conformal gravity, for instance, is a particular example of the latter. It appears as a natural ultraviolet completion of general relativity, although its symmetry becomes anomalous at one-loop level Adler (1982) (for a review of Conformal Gravity see Mannheim (2006, 2017)). In asymptotically locally AdS spacetimes, Grumiller et al. have shown that Conformal Gravity is finite without introducing boundary counterterms Grumiller et al. (2014). This theory has a well-posed variational principle for the holographic sources, i.e. the boundary metric and the partially massless response, rendering a rich scenario for studying holography Grumiller et al. (2014). Witten and Berkovits, on the other hand, showed that the supersymmetric extension of Conformal Gravity appears naturally in string theory on twistor spaces Berkovits and Witten (2004). Furthermore, it provides a natural embedding for renormalizing Einstein gravity on asymptotically locally AdS spacetimes when Neumann boundary conditions on the Fefferman-Graham expansion are imposed Maldacena (2011); Anastasiou and Olea (2016); Anastasiou et al. (2021a). Recently, it has been shown that these boundary conditions eliminate ghost instabilities that appear due to the fourth-order nature of its field equations, rendering the theory free of pathologies Hell et al. (2023).
@@
Among the different configurations of Conformal Gravity Riegert (1984); Mannheim and Kazanas (1989), gravitational instantons are particularly relevant, as they allow one to study non-perturbative effects in quantum gravity. For instance, in Ref. Strominger et al. (1984), Strominger, Horowitz, and Myers studied asymptotically flat gravitational instantons, representing contributions to the vacuum persistence amplitude between topologically inequivalent vacua. They computed the one-loop determinant in terms of the renormalization group of invariant masses and topological properties of instantons. Since the Conformal Gravity action is a quadratic form constructed out of the Weyl tensor, they show that (anti-)self-dual instantons saturate a BPS bound, given in terms of the Chern-Pontryagin index—see Eq. (15) below. More recently, extensions of the Euclidean Taub-NUT-AdS and Eguchi-Hanson metrics have been studied in Ref. Corral et al. (2021), where their conserved charges and partition function were computed. The latter are finite due to conformal invariance without invoking boundary terms. Furthermore, these configurations were shown to be inequivalent to the generalized Taub-NUT metrics studied in Refs. Iwai and Katayama (1993); Miyake (1995).
@@
Although gravitational instantons in Conformal Gravity have been studied mainly in vacuum, therefore being Bach-flat, the presence of conformal matter backreacting on these topologically nontrivial configurations remains unexplored. In order to give a step towards this direction, one must consider a matter sector that preserves the symmetries of the gravity action. In four dimensions, this holds true for the Maxwell and Yang-Mills theories. However, this is not exclusive to gauge theories: conformally-coupled scalars can also be included without spoiling conformal invariance Charmousis et al. (2009). Indeed, in Conformal Gravity theories, the presence of a scalar field that backreacts on the metric allows for the generation of hairy black hole solutions, while other configurations can be found via conformal mappings Maeda (1989). In all black hole solutions in the Einstein frame, the scalar field is coupled minimally to gravity. Applying a conformal transformation, other solutions can be obtained in the Jordan frame, which are not physically equivalent to the untransformed ones Faraoni et al. (1999).
@@
To find electrically and magnetically charged gravitational instantons, one must include backreacting Abelian gauge fields. However, when quantum effects are taken into account in any interacting gauge theory, their low-energy limit is described in terms of effective field theories Koutsoumbas et al. (2018); Diakonou et al. (1990a, b). The standard procedure for obtaining the latter is integrating out the heavy degrees of freedom in the path integral formalism. In the context of quantum electrodynamics, Euler and Heisenberg showed that quantum effects generate gauge-invariant nonlinear terms at one-loop level Heisenberg and Euler (1936). These irrelevant operators in the infrared encode quantum effects, such as vacuum polarization and photon self-interaction, to first order in the fine structure constant. Abelian gauge theories possessing higher-order terms are generically referred to as nonlinear electrodynamics, and they are characterized by strong electromagnetic fields that reduce to Maxwell’s theory in the weak field limit. For instance, the motion of particles around a magnetically charged Euler-Heisenberg black hole with scalar hair was studied in Theodosopoulos et al. (2024). Nonlinear electrodynamics exhibits novel features, including the absence of initial singularities, an upper limit on the electric field at the origin for point-like particles, and a finite self-energy of charged particles Novello et al. (2000), as shown by Born and Infeld in their seminal paper Born and Infeld (1934).
@@
In this work, we study non-conformally Einstein gravitational instantons in Conformal Gravity, both in vacuum and coupled to nonlinear fields that remain invariant under Weyl rescalings in four dimensions. In the former case, we compute the Noether-Wald charges of the Lorentzian extension of the Kerr-NUT-AdS metric with relaxed AdS boundary conditions in Conformal Gravity found in Ref. Liu and Lu (2013). After performing an analytic continuation into the Euclidean section, we obtain a curve in parametric space where the Weyl tensor becomes (anti-)self-dual. As the Conformal Gravity action is a quadratic form of the Weyl tensor, this implies that the instanton saturates a Bogomol’nyi-Prasad-Sommerfield (BPS) bound that we obtain explicitly. Then, we consider a matter sector consisting of two different types of nonlinear fields: conformally coupled scalar fields with a self-interacting potential and ModMax electrodynamics. We obtain the generalization of the Taub-NUT-AdS and Eguchi-Hanson metrics found in Ref. Corral et al. (2021), and we show that the vanishing NUT charge limit leads to a generalized Riegert metric in the presence of conformal nonlinear matter fields. We study their global properties in detail, including the partition function, topological terms, and conserved quantities.
This work is organized as follows. In Sec. II, we briefly review Conformal Gravity in four dimensions. Then, in Sec. III, we revisit the Lorentzian Kerr-NUT-AdS extension in Conformal Gravity and compute its conserved charges. In Sec. IV, we study the Euclidean continuation of this geometry, determine the curve in parameter space along which the solution becomes (anti)-self-dual and compute the corresponding on-shell Euclidean action. In Sec. V, by using the Dunajski-Tod theorem, which provides the necessary and sufficient conditions for (anti-)self-dual spaces not to be conformally related to any Einstein space, we show that the Kerr-NUT-AdS extension satisfies the conditions of this theorem. In Sec. VI, we minimally couple non-linear conformal matter to Conformal Gravity, namely conformally coupled scalar fields and ModMax non-linear electrodynamics. Section VII is devoted to the analysis of the extensions of the Taub-NUT and Eguchi-Hanson gravitational instantons within this theory. Furthermore, by taking the static limit of the former, we obtain the generalization of the Riegert metric supported by non-linear conformal matter. We analyze the global properties of these new solutions, including their on-shell Euclidean action, temperature, topological invariants, and conserved charges derived from a Gauss law. Finally, in Sec. VIII we present a summary and discussion of the main results.
@@ II Conformal gravity
$\mathcal{E}_{\mu\nu}:=8\alpha B_{\mu\nu}\,,\;\;\;\;\;\mbox{and}\;\;\;\;\;\Theta^{\mu}:=-4\alpha\left(\delta g_{\nu\sigma}C^{\sigma\mu\nu}+W^{\rho\sigma\mu\nu}\nabla_{\rho}\delta g_{\nu\sigma}\right)\,$, (5)
respectively, where $B_{\mu\nu}=\nabla^{\lambda}C_{\mu\nu\lambda}-S^{\lambda\rho}W_{\lambda\mu\nu\rho}$ is the Bach tensor and $C_{\mu\nu\lambda}=2\nabla_{[\lambda}S_{\nu]\mu}$ is the Cotton tensor. Notice that, when evaluated on Einstein-AdS spaces, say $R_{\mu\nu}=-\tfrac{3}{\ell^{2}}g_{\mu\nu}$ the Schouten tensor becomes $S_{\mu\nu}=-\tfrac{1}{2\ell^{2}}g_{\mu\nu}$, where $\ell$ is the AdS radius. One can check that the Bach and Cotton tensors vanish automatically when this condition is imposed. Therefore, all Einstein-AdS spaces are Bach-flat, and as a consequence, solutions of Conformal Gravity in vacuum. Even more: as the Bach tensor transforms covariantly under Weyl transformations, conformally-Einstein spaces will also satisfy $B_{\mu\nu}=0$ for an arbitrary conformal factor. Nevertheless, there exists a very special class of metrics, which are Bach flat, but they are not conformally related to Einstein spaces. Necessary and sufficient conditions for their existence have been explored Kozameh et al. (1985); Gover and Nurowski (2004); Dunajski and Tod (2014), and particular examples, e.g. the LeBrun metrics, have been obtained LeBrun (1991); LeBrun et al. (1994).
@@ III Kerr-NUT-AdS solution in Conformal Gravity
The construction of stationary solutions in higher-curvature theories of gravity is certainly challenging. Since Conformal Gravity has more symmetries than General Relativity, but it reduces to the latter by imposing Neumann boundary conditions on the Fefferman-Graham expansion, it is more likely to find this class of solutions in both Lorentzian and Euclidean signatures. For instance, in Ref. Mannheim and Kazanas (1991), the extension of the dyonic Kerr-Newman-AdS solution in Conformal Gravity was studied by using the Plebanski-Demianski form. Then, in Ref. Liu and Lu (2013), the dyonic Kerr-Newman-NUT-AdS was studied. Using Boyer-Lindquist coordinates, the ansatz for the line element is
@@
There are interesting limits of this metric. For instance, the Kerr-NUT-AdS solution of general relativity is also a solution of Conformal Gravity. The latter can be achieved by taking the limit $c\to 0$, which implies that $b\to 0$ as a consequence of Eq. (8). Notice that Eq. (8) implies that the limit $c/b\to 0$ is smooth. In Einstein gravity, this spacetime was originally found by Carter in Ref. Carter (1968), whose global properties were studied by Miller in Miller (1973). Here, we notice that the limit $m\to 0$ and $c\to 0$, while keeping $b\neq 0$, leads to a non-Einstein spacetime with nontrivial mass and angular momentum—see Eq. (13) below. The general solution in Eq. (7) with nonvanishing $m$, $b$, and $c$, alongside with their particular limits, possesses two Killing vectors, i.e. $\partial_{t}$ and $\partial_{\varphi}$, being stationary and axisymmetric. Additionally, it has a horizon generated by the Killing vector field $\xi=\partial_{t}+\Omega_{+}\partial_{\varphi}$, where
@@
The thermodynamic properties of this solution have been studied in Ref. Liu and Lu (2013) in the absence of the NUT charge. The presence of the latter modifies both the mass and the angular momentum, as we see next. Even more, in the Euclidean section, it allows for a nontrivial curve in parameter space where the solution becomes (anti-)self-dual. This allows us to interpret it as a gravitational instanton of Conformal Gravity that saturates a BPS bound given in terms of the Chern-Pontryagin index.
@@ III.1 Noether-Wald charges
The non-Einstein contribution to the conserved charges is encoded in the last term proportional to the Cotton tensor. Indeed, for Einstein spaces, the latter vanishes, and the Noether prepotential reduces to the topologically renormalized Einstein-AdS gravity Miskovic and Olea (2009). If $\xi$ is a globally-defined Killing vector field, conserved charges are obtained by integrating the Noether prepotential over a codimension-2 boundary integral as
@@
IV (Anti-)Self-dual instanton in Conformal Gravity
Gravitational instantons in Conformal Gravity are regular Bach-flat solutions that can be obtained from their Lorentzian counterpart by performing the analytic continuation into the Euclidean section. In particular, from Eq. (7), the (anti-)self-dual gravitation instanton can be obtained from the Wick rotation $t\to i\tau$, $n\to-i\hat{n}$, and $a\to-i\hat{a}$, where hatted quantities denote those in Euclidean signature. The absence of conical singularities renders the solution regular; this is achieved by demanding the correct periodicity of the Euclidean time and the azimuthal angle, i.e. $\tau\sim\tau+\beta_{\tau}$ and $\varphi\sim\varphi+\beta_{\varphi}$, where $\beta_{\varphi}:=-\beta_{\tau}\hat{\Omega}_{+}$ and $\beta_{\tau}=\hat{T}_{H}^{-1}$, with $\hat{\Omega}_{+}$ and $\hat{T}_{H}$ being the analytically continued angular velocity and Hawking temperature, respectively.
@@
where $\hat{b}$ is obtained by taking the Euclidean continuation of Eq. (8). The condition (14) implies that the Weyl tensor is (anti-)self dual, i.e. $W_{\mu\nu\lambda\rho}=\pm\tilde{W}_{\mu\nu\lambda\rho}$, where $\tilde{W}_{\mu\nu\lambda\rho}:=\tfrac{1}{2}\varepsilon_{\mu\nu\alpha\beta}W^{\alpha\beta}_{\lambda\rho}$ is the dual Weyl tensor and $\epsilon_{\mu\nu\lambda\rho}$ the Levi-Civita tensor. For (anti-)self-dual configurations, the BPS bound in Conformal Gravity is given by Deser and Nepomechie (1983); Strominger et al. (1984)
@@
and that of the complex projective space is $P_{1}[\mathbb{CP}^{2}]=3$, we conclude that there exists a region in the parameter space where $P_{1}[\mathcal{M}]\leq P_{1}[\mathbb{CP}^{2}]$. Therefore, within such a region, the (anti-)self-dual instanton studied here dominates the path integral, which can be obtained to first-order in the saddle-point approximation as $\ln\mathcal{Z}\approx-I_{\rm CG}$.
V Obstructions to conformally (anti-)self-dual Einstein spaces
It is well-known that the conformal symmetry of the action (1) allows one to obtain Bach-flat solutions by performing a Weyl transformation of Einstein spaces. This is because the Bach tensor is identically zero for Einstein spaces and it transforms covariantly under Weyl rescalings of the metric [cf. Eq. (3)] with conformal weight $\Delta=-2$, that is, $B_{\mu\nu}\to\bar{B}_{\mu\nu}=e^{-2\sigma(x)}B_{\mu\nu}$. Thus, if $\bar{g}_{\mu\nu}$ is conformally related to an Einstein space, it will solve the Conformal Gravity equations automatically. However, as demonstrated in Refs. Kozameh et al. (1985); Gover and Nurowski (2004); Dunajski and Tod (2014), there exist solutions of Conformal Gravity which are not conformally Einstein. Indeed, in Ref. Dunajski and Tod (2014) it was shown that a Bach-flat Riemannian metric with anti-self-dual Weyl tensor contains a Ricci-flat metric if and only if

### v2

# Nonconformally Ricci-flat instantons in Conformal Gravity with and without nonlinear matter fields

@@ Abstract
In this work, we study nonconformally Ricci-flat gravitational instantons in four-dimensional Conformal Gravity, both in vacuum and in the presence of nonlinear conformal matter. First, the one-parameter extension of the Kerr-NUT-AdS metric is analyzed. We obtain their conserved charges by using the Noether-Wald formalism. It turns out that they receive corrections from the linear modes present in Conformal Gravity, which are properly identified. Then, we perform the analytic continuation into the Euclidean section and find the curve in parameter space along which this solution becomes regular and globally (anti)self-dual. Using the Dunajski-Tod theorem, we show that the self-dual metric is not conformally Ricci-flat. Then, the backreaction of nonlinear conformal matter is considered. In particular, we find new gravitational instantons in the presence of conformally coupled scalar fields and ModMax electrodynamics. We compute the partition function and conserved charges, which turn out to be finite by virtue of the conformal invariance of the theory. As a byproduct, we also obtain a generalization of the Riegert metric dressed with nonlinear conformal matter as a particular limit of these instantons. For all cases, we analyze the global properties, the curve in parameter space where the solutions are (anti)-self-dual, and the on-shell Euclidean action, among other features.
@@ I Introduction
The anti–de Sitter/conformal field theory (AdS/CFT) correspondence has provided a powerful tool for studying strongly interacting field theories at the fixed points of the renormalization group flow, where standard perturbative techniques are no longer valid Maldacena (1998); Witten (1998); Gubser et al. (1998). At those points, quantum field theories exhibit conformal invariance, which dictates the specific form of correlator functions of the theory. In the context of gravity, theories that possess conformal symmetry display an improved ultraviolet behavior, providing a suitable scenario for studying quantum effects. Conformal gravity, for instance, is a particular example of the latter. It appears as a natural ultraviolet completion of general relativity, although its symmetry becomes anomalous at one-loop level Adler (1982) (for a review of Conformal Gravity see Mannheim (2006, 2017)). In asymptotically locally AdS spacetimes, Grumiller et al. have shown that Conformal Gravity is finite without introducing boundary counterterms Grumiller et al. (2014). This theory has a well-posed variational principle for the holographic sources, i.e. the boundary metric and the partially massless response, rendering a rich scenario for studying holography Grumiller et al. (2014). Witten and Berkovits, on the other hand, showed that the supersymmetric extension of Conformal Gravity appears naturally in string theory on twistor spaces Berkovits and Witten (2004). Furthermore, it provides a natural embedding for renormalizing Einstein gravity on asymptotically locally AdS spacetimes when Neumann boundary conditions on the Fefferman-Graham expansion are imposed Maldacena (2011); Anastasiou and Olea (2016); Anastasiou et al. (2021a). These boundary conditions eliminate ghost instabilities that appear due to the fourth-order nature of its field equations, rendering the theory free of pathologies Lu et al. (2011); Hell et al. (2023).
@@
Among the different configurations of Conformal Gravity Riegert (1984); Mannheim and Kazanas (1989), gravitational instantons are particularly relevant, as they allow one to study nonperturbative effects in quantum gravity. For instance, in Ref. Strominger et al. (1984), Strominger, Horowitz, and Myers studied asymptotically flat gravitational instantons, representing contributions to the vacuum persistence amplitude between topologically inequivalent vacua. They computed the one-loop determinant in terms of the renormalization group of invariant masses and topological properties of instantons. Since the Conformal Gravity action is a quadratic form constructed out of the Weyl tensor, they show that (anti)self-dual instantons saturate a BPS bound, given in terms of the Chern-Pontryagin index—see Eq. (15) below. More recently, extensions of the Euclidean Taub-NUT-AdS and Eguchi-Hanson metrics have been studied in Ref. Corral et al. (2021), where their conserved charges and partition function were computed. The latter are finite due to conformal invariance without invoking boundary terms. Furthermore, these configurations were shown to be inequivalent to the generalized Taub-NUT metrics studied in Refs. Iwai and Katayama (1993); Miyake (1995).
@@
Although gravitational instantons in Conformal Gravity have been studied mainly in vacuum, therefore being Bach-flat, the presence of conformal matter backreacting on these topologically nontrivial configurations remains unexplored. In order to give a step toward this direction, one must consider a matter sector that preserves the symmetries of the gravity action. In four dimensions, this holds true for the Maxwell and Yang-Mills theories. However, this is not exclusive to gauge theories: conformally-coupled scalars can also be included without spoiling conformal invariance Charmousis et al. (2009). Indeed, in Conformal Gravity theories, the presence of a scalar field that backreacts on the metric allows for the generation of hairy black hole solutions, while other configurations can be found via conformal mappings Maeda (1989). In all black hole solutions in the Einstein frame, the scalar field is coupled minimally to gravity. Applying a conformal transformation, other solutions can be obtained in the Jordan frame, which are not physically equivalent to the untransformed ones Faraoni et al. (1999).
@@
To find electrically and magnetically charged gravitational instantons, one must include backreacting Abelian gauge fields. However, when quantum effects are taken into account in any interacting gauge theory, their low-energy limit is described in terms of effective field theories Koutsoumbas et al. (2018); Diakonou et al. (1990a, b). The standard procedure for obtaining the latter is integrating out the heavy degrees of freedom in the path integral formalism. In the context of quantum electrodynamics, Euler and Heisenberg showed that quantum effects generate gauge-invariant nonlinear terms at one-loop level Heisenberg and Euler (1936). These irrelevant operators in the infrared encode quantum effects, such as vacuum polarization and photon self-interaction, to first order in the fine structure constant. Abelian gauge theories possessing higher-order terms are generically referred to as nonlinear electrodynamics, and they are characterized by strong electromagnetic fields that reduce to Maxwell’s theory in the weak field limit. For instance, the motion of particles around a magnetically charged Euler-Heisenberg black hole with scalar hair was studied in Theodosopoulos et al. (2024). Nonlinear electrodynamics exhibits novel features, including the absence of initial singularities, an upper limit on the electric field at the origin for pointlike particles, and a finite self-energy of charged particles Novello et al. (2000), as shown by Born and Infeld in their seminal paper Born and Infeld (1934).
@@
In this work, we study nonconformally Ricci-flat gravitational instantons in Conformal Gravity, both in vacuum and coupled to nonlinear fields that remain invariant under Weyl rescalings in four dimensions. In the former case, we compute the Noether-Wald charges of the Lorentzian extension of the Kerr-NUT-AdS metric with relaxed AdS boundary conditions in Conformal Gravity found in Ref. Liu and Lu (2013). After performing an analytic continuation into the Euclidean section, we obtain a curve in parametric space where the Weyl tensor becomes (anti)self-dual. As the Conformal Gravity action is a quadratic form of the Weyl tensor, this implies that the instanton saturates a Bogomol’nyi-Prasad-Sommerfield (BPS) bound that we obtain explicitly. Then, we consider a matter sector consisting of two different types of nonlinear fields: conformally coupled scalar fields with a self-interacting potential and ModMax electrodynamics. We obtain the generalization of the Taub-NUT-AdS and Eguchi-Hanson metrics found in Ref. Corral et al. (2021), and we show that the vanishing NUT charge limit leads to a generalized Riegert metric in the presence of conformal nonlinear matter fields. We study their global properties in detail, including the partition function, topological terms, and conserved quantities.
This work is organized as follows. In Sec. II, we briefly review Conformal Gravity in four dimensions. Then, in Sec. III, we revisit the Lorentzian Kerr-NUT-AdS extension in Conformal Gravity and compute its conserved charges. In Sec. IV, we study the Euclidean continuation of this geometry, determine the curve in parameter space along which the solution becomes (anti)-self-dual, and compute the corresponding on-shell Euclidean action. In Sec. V, by using the Dunajski-Tod theorem, which provides the necessary and sufficient conditions for (anti)self-dual spaces not to be conformally related to any Ricci-flat space, we show that the Kerr-NUT-AdS extension satisfies the conditions of this theorem. In Sec. VI, we minimally couple nonlinear conformal matter to Conformal Gravity, namely conformally coupled scalar fields and ModMax nonlinear electrodynamics. Section VII is devoted to the analysis of the extensions of the Taub-NUT and Eguchi-Hanson gravitational instantons within this theory. Furthermore, by taking the static limit of the former, we obtain the generalization of the Riegert metric supported by nonlinear conformal matter. We analyze the global properties of these new solutions, including their on-shell Euclidean action, temperature, topological invariants, and conserved charges derived from a Gauss law. Finally, in Sec. VIII we present a summary and discussion of the main results.
@@ II Conformal gravity
$\mathcal{E}_{\mu\nu}:=4\alpha B_{\mu\nu}\,,\;\;\;\;\;\mbox{and}\;\;\;\;\;\Theta^{\mu}:=-4\alpha\left(\delta g_{\nu\sigma}C^{\sigma\mu\nu}+W^{\rho\sigma\mu\nu}\nabla_{\rho}\delta g_{\nu\sigma}\right)\,$, (5)
respectively, where $B_{\mu\nu}=\nabla^{\lambda}C_{\mu\nu\lambda}-S^{\lambda\rho}W_{\lambda\mu\nu\rho}$ is the Bach tensor and $C_{\mu\nu\lambda}=2\nabla_{[\lambda}S_{\nu]\mu}$ is the Cotton tensor. Notice that, when evaluated on Einstein-AdS spaces, say $R_{\mu\nu}=-\tfrac{3}{\ell^{2}}g_{\mu\nu}$ the Schouten tensor becomes $S_{\mu\nu}=-\tfrac{1}{2\ell^{2}}g_{\mu\nu}$, where $\ell$ is the AdS radius. One can check that the Bach and Cotton tensors vanish automatically when this condition is imposed. Therefore, all Einstein-AdS spaces are Bach-flat, and as a consequence, solutions of Conformal Gravity in vacuum. Even more: as the Bach tensor transforms covariantly under Weyl transformations, conformally-Einstein spaces will also satisfy $B_{\mu\nu}=0$ for an arbitrary conformal factor. Nevertheless, there exists a very special class of metrics, which are Bach flat, but they are not conformally related to Einstein spaces. Necessary and sufficient conditions for their existence have been explored Kozameh et al. (1985); Gover and Nurowski (2004); Dunajski and Tod (2014) (see also et al. (2013)), and particular examples, e.g. the LeBrun metrics, have been obtained LeBrun (1991); LeBrun et al. (1994).
@@ III Kerr-NUT-AdS solution in Conformal Gravity
The construction of stationary solutions in higher-curvature theories of gravity is certainly challenging. Since Conformal Gravity has more symmetries than General Relativity, but it reduces to the latter by imposing Neumann boundary conditions on the Fefferman-Graham expansion, it is more likely to find this class of solutions in both Lorentzian and Euclidean signatures. For instance, in Ref. Mannheim and Kazanas (1991), the extension of the dyonic Kerr-Newman-AdS solution in Conformal Gravity was studied by using the Plebanski-Demianski form. [^1] See Ref. Asuncion et al. (2025) for a recent analysis in the absence of the cosmological constant and Maxwell fields. Then, in Ref. Liu and Lu (2013), the dyonic Kerr-Newman-NUT-AdS was studied. Using Boyer-Lindquist coordinates, the ansatz for the line element is
@@
There are interesting limits of this metric. For instance, the Kerr-NUT-AdS solution of general relativity is also a solution of Conformal Gravity. The latter can be achieved by taking the limit $c\to 0$, which implies that $b\to 0$ as a consequence of Eq. (8). Notice that Eq. (8) implies that the limit $c/b\to 0$ is smooth. In Einstein gravity, this spacetime was originally found by Carter in Ref. Carter (1968), whose global properties were studied by Miller in Miller (1973). Here, we notice that the limit $m\to 0$ and $c\to 0$, while keeping $b\neq 0$, leads to a non-Einstein spacetime with nontrivial mass and angular momentum—see Eq. (13) below. The general solution in Eq. (7) with nonvanishing $m$, $b$, and $c$, alongside their particular limits, possesses two Killing vectors, i.e., $\partial_{t}$ and $\partial_{\varphi}$, being stationary and axisymmetric. Additionally, it has a horizon generated by the Killing vector field $\xi=\partial_{t}+\Omega_{+}\partial_{\varphi}$, where
@@
The thermodynamic properties of this solution have been studied in Ref. Liu and Lu (2013) in the absence of the NUT charge. The presence of the latter modifies both the mass and the angular momentum, as we see next. Even more, in the Euclidean section, it allows for a nontrivial curve in parameter space where the solution becomes (anti)self-dual. This allows us to interpret it as a gravitational instanton of Conformal Gravity that saturates a BPS bound given in terms of the Chern-Pontryagin index.
@@ III.1 Noether-Wald charges
The non-Einstein contribution to the conserved charges is encoded in the last term proportional to the Cotton tensor. Indeed, for Einstein spaces, the latter vanishes, and the Noether prepotential reduces to the topologically renormalized Einstein-AdS gravity Aros et al. (2000); Olea (2007); Miskovic and Olea (2009), also known as Kounterterms. If $\xi$ is a globally defined Killing vector field, conserved charges are obtained by integrating the Noether prepotential over a codimension-2 boundary integral as
@@
IV (Anti)Self-dual instanton in Conformal Gravity
Gravitational instantons in Conformal Gravity are regular Bach-flat solutions that can be obtained from their Lorentzian counterpart by performing the analytic continuation into the Euclidean section. In particular, from Eq. (7), the (anti)self-dual gravitation instanton can be obtained from the Wick rotation $t\to i\tau$, $n\to-i\hat{n}$, and $a\to-i\hat{a}$, where hatted quantities denote those in Euclidean signature. The absence of conical singularities renders the solution regular; this is achieved by demanding the correct periodicity of the Euclidean time and the azimuthal angle, i.e. $\tau\sim\tau+\beta_{\tau}$ and $\varphi\sim\varphi+\beta_{\varphi}$, where $\beta_{\varphi}:=-\beta_{\tau}\hat{\Omega}_{+}$ and $\beta_{\tau}=\hat{T}_{H}^{-1}$, with $\hat{\Omega}_{+}$ and $\hat{T}_{H}$ being the analytically continued angular velocity and Hawking temperature, respectively.
@@
where $\hat{b}$ is obtained by taking the Euclidean continuation of Eq. (8). The condition (14) implies that the Weyl tensor is (anti-)self dual, i.e. $W_{\mu\nu\lambda\rho}=\pm\tilde{W}_{\mu\nu\lambda\rho}$, where $\tilde{W}_{\mu\nu\lambda\rho}:=\tfrac{1}{2}\varepsilon_{\mu\nu\alpha\beta}W^{\alpha\beta}_{\lambda\rho}$ is the dual Weyl tensor and $\epsilon_{\mu\nu\lambda\rho}$ the Levi-Civita tensor. For (anti)self-dual configurations, the BPS bound in Conformal Gravity is given by Deser and Nepomechie (1983); Strominger et al. (1984)
@@
and that of the complex projective space is $P_{1}[\mathbb{CP}^{2}]=3$, we conclude that, since the parametric space of this (anti)-self dual instanton is two-dimensional, there is always a region in the parameter space such that the condition $P_{1}[\mathcal{M}]\leq P_{1}[\mathbb{CP}^{2}]$ is satisfied. Therefore, within such a region, the (anti)self-dual instanton studied here would dominate the path integral if the one-loop determinant is not imaginary. Indeed, one can use the techniques developed in Ref. Strominger et al. (1984) for gravitational instantons in Conformal Gravity to prove this. We left this proof for future work.
V Obstructions to conformally (anti)self-dual Ricci-flat spaces
It is well-known that the conformal symmetry of the action (1) allows one to obtain Bach-flat solutions by performing a Weyl transformation of Einstein spaces. This is because the Bach tensor is identically zero for Einstein spaces, and it transforms covariantly under Weyl rescalings of the metric [cf. Eq. (3)] with conformal weight $\Delta=-2$, that is, $B_{\mu\nu}\to\bar{B}_{\mu\nu}=e^{-2\sigma(x)}B_{\mu\nu}$. Thus, if $\bar{g}_{\mu\nu}$ is conformally related to an Einstein space, it will solve the Conformal Gravity equations automatically. However, as demonstrated in Refs. Kozameh et al. (1985); Gover and Nurowski (2004); Dunajski and Tod (2014), there exist solutions of Conformal Gravity which are not conformally Ricci-flat. Indeed, in Ref. Dunajski and Tod (2014) it was shown that a Bach-flat Riemannian metric with anti-self-dual Weyl tensor contains a Ricci-flat metric within its conformal class if and only if